### v2

# Many-body critical phase in a quasiperiodic chain and dynamical Widom lines in Fock space properties

Abstract
We study a quasiperiodic model in one dimension, namely the extended Aubry-André-Harper (EAAH) chain, that realizes a critical phase comprising entirely single-particle critical states in the non-interacting limit. In the presence of short-range interactions, the non-interacting critical phase transforms to a many-body critical (MBC) phase, separated by lines of MBC-ergodic, MBC-many-body localized (MBL) and ergodic-MBL phase transitions that meet at a triple point. We elucidate the unusual characteristics of the MBC phase compared to the ergodic and MBL phases through the localization properties of the excitations in real space and Fock space (FS), and eigenstate inverse participation ratio (IPR). We show that the MBC phase, like the MBL phase, is well described by a multifractal scaling of the IPR and a linear finite-size scaling ansatz near the transition to the ergodic and MBL phases. However, the MBC phase, at the same time, exhibits delocalization of all single-particle excitations and a system-size dependent Fock-space localization length, analogous to the ergodic phase. Remarkably, we find evidence of unusual Widom lines on the phase diagram in the form of lines of pronounced peaks or dips in the FS localization properties inside the MBC and MBL phases. These Widom lines either emerge as a continuation of the precursor phase transition line, terminating at the triple point, or originate from a phase boundary.
I Introduction
Following Anderson’s seminal work Anderson (1958), single-particle eigenstates of a non-interacting quantum particle are known to be either localized or delocalized in systems with quenched randomness, e.g., systems with random disorder potentials, due to the phenomenon of Anderson localization (AL). The AL phenomenon can even occur in systems with quasiperiodic potentials, as exemplified by the the well-known Aubry-André-Harper (AAH) model Aubry and André (1980). However, a third category of states, namely critical states, intermediate to localized and delocalized states, are known to exist right at the critical point corresponding to the Anderson delocalization-localization transition. The Anderson transition can be realized in random systems in dimensions greater than one, like in 2D or 3D depending on symmetries Abrahams et al. (1979); Evers and Mirlin (2008), and even in 1D for quasiperiodic systems Aubry and André (1980). At the Anderson transition, all the eigenstates turn critical showing remarkable multifractal properties Evers and Mirlin (2008). An outstanding question is whether one can realize a critical phase, hosting entirely critical states over an extended region of parameter space, instead of at a single critical point. Moreover, the existence of such a critical phase leads to a further natural question of the stability of the phase in the presence of many-body interactions.
Indeed there are non-interacting models, albeit very limited in number, with extended critical phases. All these models have quasiperiodic potentials. However, in addition, one needs to introduce either quasiperiodic hopping  Han et al. (1994); Chang et al. (1997); Liu et al. (2016), a Fibonacci-sequenced potential Kohmoto et al. (1987); Lin et al. (2023), spin-orbit coupling Wang et al. (2022) or $p$-wave pairing Wang et al. (2016a, b) alongside a quasiperiodic potential. Some of these models have also been experimentally realized recently using superconducting circuits Li et al. (2023) and ultra-cold atoms Xiao et al. (2021); Wang et al. (2020). There also exist a few one-dimensional models that show a phase with a fraction of critical states in the eigenspectrum Zhou et al. (2022); Deng et al. (2019); Roy and Sharma (2021); Lee et al. (2023); Ahmed et al. (2022). Once interactions are incorporated in the above systems, one expects the critical phase to be much more fragile compared to the Anderson localized phase with all single-particle states localized. In the last two decades, the stability of the Anderson localized phase in the presence of interactions, leading to the phenomenon of many-body localization (MBL) Abanin and Papić (2017); Alet and Laflorencie (2018); Abanin et al. (2019), has been indicated by many theoretical works Basko et al. (2006); Gornyi et al. (2005); Oganesyan and Huse (2007); Pal and Huse (2010); Luitz et al. (2015); Imbrie (2016); De Roeck and Huveneers (2017), at least, in 1D. Evidence of MBL have also been found in experiments Schreiber et al. (2015); Rispoli et al. (2019); Lukin et al. (2019), although the regime of stability Sels and Polkovnikov (2021); Morningstar et al. (2022); Crowley and Chandran (2022); Long et al. (2023); Sels (2022); Šuntajs et al. (2020); Sierant and Zakrzewski (2022) of the MBL phase still remains under active debate. The MBL phase exhibits many fascinating properties, such as a lack of thermalization or ergodicity Alet and Laflorencie (2018); Abanin et al. (2019); Sierant et al. and a violation of the eigenstate thermalization hypothesis (ETH) Deutsch (1991); Srednicki (1994) in isolated quantum systems. Universal properties of the putative thermal-MBL critical point have been studied in numerous works Luitz et al. (2015); Zhang and Yao (2018); Dumitrescu et al. (2019); Kiefer-Emmanouilidis et al. (2020); Abanin et al. (2021); Sierant et al. (2020); Laflorencie et al. (2020); Kudo and Deguchi (2018); Bertrand and García-García (2016); Setiawan et al. (2017); Sutradhar et al. (2022); Vosk and Altman (2013); Goremykina et al. (2019); Morningstar et al. (2020); De Roeck and Huveneers (2017); Morningstar et al. (2020).
Intriguingly, a recent numerical work Wang et al. (2021) has shown the emergence of a many-body critical phase (MBC), as opposed to a thermal-MBL critical point, out of the non-interacting critical phase, when interactions are introduced in a fermionic extended AAH (EAAH) model with quasiperiodic hopping Han et al. (1994). Ref.[Wang et al., 2021] classifies the eigenstates of the MBC phase as nonergodic extended (NEE) based on sub-thermal volume-law scaling of entanglement entropy of the states and the critical level statistics in the phase. However, when viewed from the perspective of a Fock space (FS) lattice or graph Welsh and Logan (2018); Logan and Welsh (2019); Roy and Logan (2020a, 2021); Roy (2022); Roy et al. (2023); Ghosh et al. (2019, 2025), many-body eigenstates in the MBL phase, or even in the non-interacting Anderson localized phase, can be characterized as non-ergodic extended based on the fractal or multi-fractal scaling of the inverse participation ratio (IPR) Roy and Logan (2020a, 2021) or the local FS propagator Sutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025) with the number of FS basis states $\mathcal{N}_{F}$. This situation is different from random regular graphs with uncorrelated disorder where the existence of NEE states has been highly debated Altshuler et al. (2016); Tikhonov et al. (2016); Kravtsov et al. (2018); Herre et al. (2023); Vanoni et al. (2024). Yet, another type of NEE state Li et al. (2015); Modak and Mukerjee (2015); Deng et al. (2017); Modak et al. (2018); Ghosh et al. (2020) has been found in systems with a particular type of quasiperiodic disorder, namely the generalized Aubry-André-Harper (GAAH) model  Ganeshan et al. (2015); Modak and Mukerjee (2015). These NEE states in the interacting GAAH model presumably arise due to the existence of a single-particle (SP) mobility edge in the non-interacting limit of the GAAH model Modak and Mukerjee (2015); Ghosh et al. (2019); Roy et al. (2023), and the interaction-induced coupling between localized and delocalized single-particle states across the SP mobility edge.
In this work, we ask how the NEE states in the MBC and non-interacting critical phases of the EAAH model are different in terms of their Fock and real-space properties from non-interacting delocalized and localized, ergodic, MBL and other previously studied NEE states Modak and Mukerjee (2015); Ghosh et al. (2020); Deng et al. (2017); Roy et al. (2023). We also study the nature of the transitions from delocalized and localized phases to the critical phase in the non-interacting model, and the transitions from the ergodic and MBL phases to the MBC phase in the interacting model. Intriguingly, we find the evidence of unusual Widom-like lines Xu et al. (2005); Franzese and Stanley (2007); Simeoni et al. (2010); Luo et al. (2014); Sordi and Tremblay (2024) on the dynamical phase diagram of the EAAH model, namely the existence of loci of pronounced peaks or dips in the FS localization properties deep inside the MBC (critical) and MBL (localized) phases. These Widom-like lines either appear as a continuation of the antecedent phase transition line or emanate from a phase boundary.
To this end, we compute the local density of states (LDOS) for single-particle excitations, to characterize the many-body eigenstates in the delocalized, critical and localized phases of the non-interacting EAAH model, and ergodic, MBC and MBL phases of the interacting EAAH model. The LDOS captures localization properties of the single-particle excitations in real space. To obtain a complementary perspective, the localization properties in these phases are quantified in FS via IPR, and the local and non-local FS propagators Sutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025). A FS lattice or graph is constructed using the many-body basis of occupations of the real-space sites. Any non-interacting or interacting model on a real-space lattice can be viewed as a non-interacting hopping problem on the FS lattice Altland and Micklitz (2017); Welsh and Logan (2018); Logan and Welsh (2019); Roy and Logan (2020a, 2021); Ghosh et al. (2019); Sutradhar et al. (2022). This approach has been explored in many recent works to emphasize the role of correlated disorder in the FS to give rise to MBL phenomena Pietracaprina et al. (2017); Logan and Welsh (2019); Altland and Micklitz (2017); Welsh and Logan (2018); Roy and Logan (2020a, 2021); Ghosh et al. (2019); Scoquart et al. (2024) and the inevitable multifractal nature of the MBL eigenstates De Luca and Scardicchio (2013); Luitz et al. (2015); Macé et al. (2019); Roy and Logan (2021); De Tomasi et al. (2021). Subsequently, numerical scaling theories of the MBL transition in terms of the FS inverse participation ratio (IPR) and FS propagator  Macé et al. (2019); Roy and Logan (2021); Sutradhar et al. (2022) have been developed based on the FS perspective. The latter has been useful Roy et al. (2023) to distinguish thermal and MBL states from the NEE states in the GAAH model.
In FS, we look into the scaling of IPR with FS dimension or the number of sites on the FS lattice $\mathcal{N}_{F}$, i.e., $IPR\simeq A_{I}\mathcal{N}_{F}^{-D_{2}}$, and the scaling of the typical value of the imaginary part of the local FS self energy, $\Delta_{t}\simeq A_{s}\mathcal{N}_{F}^{-(1-D_{s})}$. Here $D_{2}$ and $D_{s}$ are fractal and spectral fractal dimensions, respectively, and $A_{I}$ and $A_{s}$ are the amplitudes of the power laws. One expects, $D_{2}=D_{s}=1$, $D_{2}=D_{s}=0$ and $0<D_{s}<1$, respectively, for the delocalized, localized and non-ergodic extended states in FS. Our main results for the EAAH model, summarized in Table 1, are the following.
PHASES DIAGNOSTICS
Single particle (SP) excitations in real-space FS Inverse participation ratio $(IPR=A_{I}\mathcal{N}_{F}^{-D_{2}})$ Typical value of FS self-energy $(\Delta_{t}=A_{s}\mathcal{N}_{F}^{-(1-D_{s})})$ FS localization length $(\xi_{F})$
Delocalized All SP excitations delocalized $A_{I}>1$, $D_{2}=1$ $A_{s}<1$, $D_{s}=1$ $\xi_{F}$ increases with $L$ (slower than ergodic)
Critical All SP excitations delocalized $A_{I}>1$, $0<D_{2}<1$ Anomalous $L$-dependence $\xi_{F}$ increases with $L$ (slower than MBC)
Localized All SP excitations localized $A_{I}<1$, $0<D_{2}<1$ $A_{s}>1$, $0<D_{s}<1$ (deep within the phase) $\xi_{F}$ independent of $L$ (more robust than MBL)
Ergodic All SP excitations delocalized $A_{I}>1$, $D_{2}=1$ $A_{s}<1$, $D_{s}=1$ $\xi_{F}$ increases with $L$
MBC All SP excitations tend to delocalize for $L\rightarrow\infty$ $A_{I}>1$, $0<D_{2}<1$ Anomalous $L$-dependence $\xi_{F}$ increases with $L$
MBL All SP excitations localized $A_{I}<1$, $0<D_{2}<1$ $A_{s}>1$, $0<D_{s}<1$ $\xi_{F}$ independent of $L$
Table 1: Classification of many-body phases of the noninteracting (delocalized, critical and localized) and interacting (ergodic, MBC and MBL) systems: based on various real-space and Fock-space diagnostics. Here $L$ and $\mathcal{N}_{F}$ are the number of sites on the real-space and Fock-space lattices, respectively
(1) By calculating the typical value $\rho_{t}(\omega)$ of the LDOS at a SP excitation energy $\omega$, we find that all SP excitations in real-space are delocalized in the non-interacting critical and MBC phases. This feature of the critical phases is like that in the delocalized and ergodic phases, but unlike in the localized and MBL phases. The delocalized nature of all SP excitations in the critical phases is in contrast to that of the NEE phase in the GAAH model Roy et al. (2023), where there exists a mobility edge between localized and delocalized SP excitations.
(2) From the scaling of IPR, we find that critical and MBC states are indeed extended but non-ergodic, having non-zero amplitude over an infinite number, $\mathcal{N}_{F}^{D_{2}}\to\infty$ ($0<D_{2}<1$), albeit zero fraction of sites in the thermodynamic limit $\mathcal{N}_{F}\to\infty$. Thus, in this regard, the critical phases are similar to the multifractal states with $0<D_{2}<1$ in the localized and MBL phases, as well as the NEE phase Roy et al. (2023) in the GAAH model, and dissimilar to the delocalized and ergodic states, where $D_{2}=1$.
(3) We find the existence of Widom-like lines in the form of a line of peaks or dips in FS quantities like $IPR$ and $\Delta_{t}$ inside the MBC (critical) and MBL (localized) phases. Such a line of non-monotonicity in physical quantities inside a phase is reminiscent of the well-known Widom or Fisher-Widom lines obseved in a supercritical fluid Xu et al. (2005); Franzese and Stanley (2007); Simeoni et al. (2010); Luo et al. (2014); Sordi and Tremblay (2024) in the thermodynamic phase diagram. Here, however, we detect these lines on the dynamical phase diagram of the EAAH model.
(4) We find that the typical self energy $\Delta_{t}(\mathcal{N}_{F})$ does not follow any systematic scaling with $\mathcal{N}_{F}$ in the non-interacting critical and MBC phases, often exhibiting a non-monotonic dependence with $\mathcal{N}_{F}$, as well as with the parameters of the EAAH model. This is in contrast to the well-defined scaling behaviour $\Delta_{t}\sim\mathcal{N}_{F}^{-(1-D_{s})}$, seen in the delocalized and ergodic phases with $D_{s}=1$, AL and MBL phases with $0<D_{s}<1$, and the NEE phaseRoy et al. (2023) in the GAAH model with $0<D_{s}<1$.
(5) For the finite system sizes accessible by our ED calculations, we find that the ergodic-MBC transition is well described by the same asymmetric scaling ansatz Garcia-Mata et al. (2017) that has been found Macé et al. (2019); Roy and Logan (2021); Sutradhar et al. (2022) to describe the ergodic-MBL transition for similar system sizes. In particular, we find a volumic finite-size scaling form, governed by a nonergodic FS volume which diverges with a Kosterlitz-Thouless (KT) like essential singularity for both the ergodic-MBL and ergodic-MBC transitions in the EAAH model. On the other hand, a linear scaling form with power-law divergent correlation length describes the transitions approaching from the MBC and MBL phases.
(6) We extract a FS length scale $\xi_{F}$ from the decay of non-local FS propagators on the FS lattice. $\xi_{F}$ increases with $L$ in the MBC (critical) phase as in the ergodic (delocalized) phase and NEE phase Roy et al. (2023) in the GAAH model. This is in contrast to MBL (AL) phase where $\xi_{F}$ is almost independent of $L$. We find that $\xi_{F}(L)$ increases much faster with $L$ in the interacting ergodic and MBC phases compared to the non-interacting delocalized and critical phases.
As shown in Table 1, the amplitude $A_{I}$ of the power-law scaling of IPR in ergodic (delocalized) and MBC (critical) phases is greater than 1, whereas in the MBL (AL) phase $A_{I}<1$. The ergodic, MBC and MBL phases on the FS qualitatively look similar to the corresponding non-interacting phases in terms of the typical LDOS, FS IPR and self-energy. However, the distribution of the IPR and self energy show quite different behaviors in the ergodic phase and non-interacting delocalized phase.
The rest of the paper is organized as follows. In Sec. II, we describe the model. In Sec. III we define the single-particle excitations in real space and then use them for characterization of the phases for noninteracting and interacting systems, in Secs. III.1 and III.2, respectively. Then in Sec. IV, we define the FS propagator and discuss the FS lattice structure. In Secs. IV.1 and IV.2, we calculate the imaginary part of the Feenberg self energy and inverse participation ratio in FS to characterize phases in noninteracting and interacting systems respectively followed by a finite size scaling of inverse participation ratio across different phase transitions in the interacting systems. In Sec. V, we define the FS localization length and distinguish different phases using it. Then in Sec. VI we discuss possibility of getting the Widom-like line in presence of random potential. Finally, we conclude in Sec. VII.
II Model
The (EAAH) model Hamiltonian of the interacting spinless fermions is given by Wang et al. (2021),
$H=\sum_{i=1}^{L-1}(t_{i}c_{i}^{\dagger}c_{i+1}+h.c.)+\sum_{i=1}^{L}h_{i}n_{i}+V\sum\limits_{i=1}^{L-1}n_{i}n_{i+1}$ (1)
     Refer to caption(a)           Refer to caption(b)
Figure 1: Phase diagrams: (a) Phase diagram of the noninteracting $(V=0)$ EAAH model. The phases are delocalized, localized and critical separated by the solid lines. The vertical dashed line ($\lambda=2$) in the critical phase and horizontal dashed line ($\mu=1$) in the localized phase indicate the non-monotonic behavior (‘Widom line’) in IPR and $\Delta_{t}$ which are also used to obtain the phase boundaries. (b) Similar phase diagram for interacting $(V=1)$ EAAH model with ergodic, MBL and MBC phases separated by solid lines. The dashed lines in MBC and MBL phases denote similar Widom lines, implying non-monotonic behavior in IPR and $\Delta_{t}$ inside the phases, respectively. Here the strengths of the onsite quasi-periodic potential and quasi-periodic hoppings are denoted by $\lambda$ and $\mu$, respectively.
Here $c_{i}$ is the fermion annihilation operator at site $i$ and $n_{i}=c^{\dagger}_{i}c_{i}$. We study the model at half filling. The nearest-neighbor hopping strength $t_{i}=J+\mu\cos(2\pi\chi(i+1/2)+\phi)$ and the nearest-neighbor interaction strength $V=1$. The quasiperiodic potential $h_{i}=\lambda\cos(2\pi\chi i+\phi)$, where $\chi=(\sqrt{5}-1)/2$ and $\lambda$ is the strength of potential with a global phase $\phi\in(0,2\pi]$. The phase $\phi$ is used to generate different realizations of disorder and perform disorder averaging. By taking $\phi=0$ without loss of generality, followed by a special kind of Fourier transformation Aubry and André (1980); Biddle and Das Sarma (2010) $c_{j}=\frac{1}{\sqrt{L}}\sum_{k}e^{-i2\pi jk\chi}c_{k}$, Eq. 1 for $V=0$ in a closed chain with periodic boundary condition yields
$H=\sum_{k}(t^{{}^{\prime}}_{k}c_{k}^{\dagger}c_{k+1}+h.c.)+\sum_{k}h^{{}^{\prime}}_{k}c_{k}^{\dagger}c_{k},$ (2)
where $t^{{}^{\prime}}_{k}=\lambda/2+\mu\cos(2\pi\chi(k+1/2))$ and $h^{{}^{\prime}}_{k}=2J\cos(2\pi\chi k)$ . Hence the Hamiltonians in Eq. 2 and Eq. 1 in the non-interacting limit ($V=0$) are related by a duality transformation and the model is self-dual at $\lambda=2J$ . For $\mu=0$, one obtains the Aubry-André-Harper (AAH) model, with the self duality point $\lambda=2J$, which coincides with the delocalization-localization transition Aubry and André (1980). The above consideration suggests that the same AAH self-duality also exists for $\mu\neq 0$, and $\lambda=2J$ remains the delocalization to localization phase transition. This is indeed the case as shown in Fig.1(a). Moreover, the model exhibits further intriguing features for $\mu\geq J$. In this case, a distribution of zeros appears in the quasiperiodic hopping $t_{i}$’s in the thermodynamic limit such that the system can no longer support a delocalized eigenstate. This mechanism has been put forward recently as a way to produce the critical states Zhou et al. (2022). Only critical or localized eigenstates can exist in this parameter region. As a result, for small values of onsite potential strength $\lambda$ all the delocalized eigenstates turn out to be critical with $\mu=J$ being the delocalized-critical transition line for $\lambda<2J$. For $\lambda>2J$, one obtains another critical line $\lambda=2\mu$ separating the critical and localized phases, which was obtained numerically Han et al. (1994); Chang et al. (1997); Takada et al. (2004); Aditya and Roy (2024). We set $J=1$ in our calculations. The schematic of the phase diagram of the non-interacting EAAH model is shown in Fig. 1(a). The phase diagram has been obtained numerically based on earlier works Han et al. (1994); Chang et al. (1997); Takada et al. (2004); Aditya and Roy (2024) and various diagnostics used in this work, as discussed later.
Given the real-momentum space duality of the non-interacting model under $\lambda\rightleftarrows 2J$ for any $\mu$, we find that $\lambda=2$ remains to be the self-dual line, even in the critical phase for $\mu>1$. However, the phase boundaries $\lambda=2\mu$ for $\lambda>2$ and $\mu=1$ for $\lambda<2$ in Fig.1(a) are dual to one another. As we discuss later, Widom like lines $\lambda=2,\mu>1$ and $\mu=1,\lambda>2$ (dual to $\lambda=2\mu$, $\mu<1$ in delocalized phase) appear in the phase diagram for the non-interacting model, both in real-space and Fock-space quantities. Remarkably, the vestiges of these lines seem to persist as Widom-like lines even in the interacting model ($V\neq 0$) where the self-duality is broken. In principle, one may expect to find a $\lambda=2\mu$ line for $\mu<1$ (dual to $\mu=1$ line for $\lambda>2$ in localized phase) in the non-interacting delocalized phase. Nevertheless, no strong signature of this line is seen within the delocalized phase for the quantities studied in this work.
In the presence of repulsive many-body nearest-neighbor interactions, the self-duality is broken. As a result, all the critical phase boundaries shift in parameter space except the ergodic-MBC transition that still appears at $\mu=1$, as shown in the schematic Fig.1(b). This is because the argument based on the (near) zeros of the quasiperiodic hopping, as explained previously, still holds in the interacting limit where the interaction strength $V$ presumably acts as an irrelevant parameter. The many-body phases and critical lines have been obtained in Ref. [Wang et al., 2021] by calculating mainly the energy-level statistics and the half-chain entanglement entropy of the half-filled system. The level spacing ratio $\langle r\rangle=0.53$ and $0.386$ in the ergodic and MBL phases, respectively whereas $\langle r\rangle$ shows an intermediate value in the MBC phase. A further analysis of the many-body energy level spacings and their correlations can be found in Appendix A. In the ergodic phase, the half-chain eigenstate entanglement entropy $S_{A}$ shows the thermal volume-law scaling ($\propto L$), with a coefficient of proportionality consistent with the thermal entropy density in the thermodynamic limit Wang et al. (2021), and $S_{A}=S_{A}^{th}$. In the MBL phase, $S_{A}$ follows area-law scaling, whereas in the MBC phase, the entanglement entropy is sub-thermal, $S_{A}<S_{A}^{Th}$, although $S_{A}$ still satisfies a volume law. A schematic of the phase diagram the of the interacting system is shown in Fig. 1(b) based on earlier works Han et al. (1994); Chang et al. (1997); Takada et al. (2004); Aditya and Roy (2024) and our results.
III Real-space single-particle excitations
[…]
In particular, we calculate the spectral function characterized by local density of states (LDOS) $\rho_{n}(i,\omega)=-(1/\pi)Im[\mathcal{G}_{n}(i,\omega)]$ for an excitation energy $\omega$ at the middle of the many-body energy spectrum ($E_{n}\simeq 0$) for the interacting EAAH model. The LDOS does not depend on many-body energy for the non-interacting case ($V=0$). We obtain the typical LDOS defined as $\ln\rho_{t}(\omega)=\langle\ln\rho_{n}(i,\omega)\rangle$ and average LDOS defined as $\rho_{a}(\omega)=\langle\rho_{n}(i,\omega)\rangle$, where $\langle$$\rangle$ denotes an arithmetic average over the lattice sites and $\phi$. The single particle excitations originate from the poles of the Green’s function $\mathcal{G}_{n}(i,\omega)$. For a localized state, a finite number of discrete poles, corresponding to the localization volume (length), gives rise to discrete peaks of $\mathcal{G}_{n}(i,\omega)$ in $\omega$ having zero measure in the LDOS even in the thermodynamic limit. In contrast, for a delocalized state the poles of $\mathcal{G}_{n}(i,\omega)$ form a continuum for $\omega$ lying within the energy band. Hence, $\rho_{t}(\omega)\rightarrow 0$ and $\mathcal{O}(1)$ in the thermodynamic limit for localized and delocalized phases, respectively Ganeshan et al. (2015). On the other hand, the arithmetic mean $\rho_{a}(\omega)$ approaches a non-zero value with increasing system size both in the localized and delocalized phases. Thus $\rho_{t}(\omega)$ or the ratio $\rho_{t}(\omega)/\rho_{a}(\omega)$ can be used as an order parameter to detect the localization of an excitation with energy $\omega$ Dobrosavljević et al. (2003); Dobrosavljević and Kotliar (1997); Jana et al. (2021); Roy et al. (2023).
III.1 Noninteracting systems
In the noninteracting limit $(V=0)$, Eq. 3 can be simplified to write the LDOS as,
$\rho(i,\omega)=\frac{1}{\pi}\sum\limits_{k}|\psi_{k}(i)|^{2}\frac{\eta}{(\omega-\epsilon_{k})^{2}+\eta^{2}},$ (4)
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 2: Real-space single-particle excitations in noninteracting system: (a-b) Typical values of LDOS $\rho_{t}(\omega)$ for increasing $L$ in the delocalized and localized phases, for $\mu=0.5,~\lambda=1.0$ and $\mu=0.5,~\lambda=5.0$, respectively. The dashed curves in the figure represent the gapped excitations, where both $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ (not shown) decrease with $L$. (c-d) $\rho_{t}(\omega)$ for increasing $L$ at two different points, $(\mu=1.5,\lambda=1.0)$ and $(\mu=2.0,\lambda=1.0)$, in the critical phase on the phase diagram of Fig.1(a).
where $\psi_{k}(i)$ and $\epsilon_{k}$ are the single particle eigenfunction and energy, respectively. The $L$-dependence of both $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ can be useful to detect different types of SP excitations, delocalized, localized and gapped. Gaps are $\mathcal{O}(1)$ interval of $\omega$ in the single particle energy spectra without any spectral weight in the thermodynamic limit $L\to\infty$. Eq. 4 can be further written as $\rho(i,\omega)\approx(1/\pi\eta)\sum_{|\omega-\epsilon_{k}|\lesssim\eta}|\psi_{k}(i)|^{2}+(\eta/\pi)\sum_{|\omega-\epsilon_{k}|\gtrsim\eta}|\psi_{k}(i)|^{2}/(\omega-\epsilon_{k})^{2}$ . For an excitation energy $\omega$ that falls within a gap, the second term contributes and $\rho(i,\omega)\sim|\psi_{k}(i)|^{2}/L$, where $\psi_{k}$ corresponds to state(s) at the gap edge withing the band. This leads to both $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ decreasing at least as fast as $1/L$. For excitation energy $\omega$ within an energy band, $\rho(i,\omega)\sim L|\psi_{k}(i)|^{2}+\mathcal{O}(1/L)$. Now, for delocalized and localized states $|\psi_{k}(i)|^{2}\sim 1/L$ and $|\psi_{k}(i)|^{2}\sim e^{-|i-i_{k}|/\xi}$, respectively, where $\xi$ is the single particle localization length and $i_{k}$ is localization center. As a result, for delocalized SP excitations, $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ are similar and remains nonzero in the thermodynamic limit. In contrast, the typical value of $\rho(i,\omega)$, $\rho_{t}(\omega)$, vanishes for the localized SP excitations for $L\to\infty$. However, the arithmetic mean $\rho_{a}(\omega)$ remains nonzero in the thermodynamic limit due to averaging over $i$ and/or disorder realizations, since the discrete peaks of $\rho(i,\omega)$ form a continuum when sampled over all sites and/or disorder realizations.
The results for $\rho_{t}(\omega)$ are shown in Fig. 2(a) and Fig. 2(b) for delocalized and localized phases of the EAAH model, where all the SP excitations are delocalized and localized, respectively. The dashed curves in the figure represent the gapped excitations, where both $\rho_{t}(\omega)$ and $\rho_{a}(\omega)$ (not shown) decrease with $L$. Interestingly, in the critical phase $\rho_{t}(\omega)$ saturates to $L$-independent nonzero values, as shown in Figs. 2(c) and 2(d) for two different points in the critical phase, qualitatively similar to that of the delocalized phase.
III.2 Interacting systems
In this subsection, we characterize the phases of the interacting EAAH model through the spectral function, namely, the typical value $\rho_{t}(\omega)$ of LDOS for many-body eigenstates chosen from a sufficiently small energy window at the middle of the many-body energy spectrum. In particular, we ask how the SP excitations in the delocalized, localized and critical phases of the non-interacting EAAH model, discussed in the preceding section, are impacted by the effects of interaction in the ergodic, MBL and MBC phases of the interacting model. To obtain $\rho_{t}(\omega)$, We perform an average over $2000,500,250$ and $100$ samples of global phase $\phi$ for system of sizes $L=10,12,14$ and $16$, respectively. Typical LDOS $\rho_{t}(\omega)$ for excitation of energy $\omega$ in the ergodic and MBL phases are shown in Fig. 3(a) and Fig. 3(b), respectively. In the ergodic phase, $\rho_{t}(\omega)$ saturates to a finite value with increasing $L$, indicating delocalization of all excitations within the bandwidth $|\omega|\lesssim 4$. In contrast, in the MBL phase $\rho_{t}(\omega)$ decreases rapidly with $L$. For example, we have verified that $\rho_{t}(\omega)/\rho_{a}(\omega)\sim e^{-L}$. This implies localization of SP excitations at all $\omega$, as also found for the GAAH model Roy et al. (2023). In Figs. 3(c-d), we show $\rho_{t}(\omega)$ for two different choices of parameters in the MBC phase. Qualitatively, in both the figures, $\rho_{t}(\omega)$ tends to saturate to a finite value as $L$ increases, albeit with more pronounced $L$ dependence in compared to that for the ergodic phase in Fig. 3(a). The saturation of $\rho_{t}(\omega)$ to a finite value indicates delocalization of the single particle exitations at all energies $\omega$. However, the saturation value of $\rho_{t}(\omega)$ in the MBC phase is order of magnitude smaller than that in the ergodic phase. This is also seen by comparing (not shown) the typical LDOS $\rho_{t}(\omega)$ with mean LDOS $\rho_{a}(\omega)$ in these phases. In the the ergodic phase, $\rho_{t}(\omega)\lesssim\rho_{a}(\omega)$, whereas $\rho_{t}(\omega)\ll\rho_{a}(\omega)$ in the MBC phase. In contrast, in the non-interacting EAAH model, $\rho_{t}(\omega)\lesssim\rho_{a}(\omega)$ for both delocalized and critical phases.
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 3: Real-space single-particle excitations in interacting system: (a-b) Typical values of LDOS $\rho_{t}(\omega)$ as a function of system size $L$ in ergodic and MBL phases, for $\mu=0.5,~\lambda=2.0$ and $\mu=0.5,~\lambda=5.0$, respectively. (c-d) $\rho_{t}(\omega)$ for increasing $L$ at two different points, $(\mu=1.5,\lambda=1.0)$ and $(\mu=2.0,\lambda=1.0)$, in the MBC phase on the phase diagram of Fig.1(b).
$\rho_{t}(\omega)$ in the MBC phase of interacting systems shows qualitatively similar behavior as that in the corresponding critical phase of the noninteracting systems, although $L$-dependence for the non-interacting critical phase in Fig.2(c) cannot be detected due to much larger system sizes accessed there for the non-interacting EAAH. These results are in contrast with that in an earlier work Roy et al. (2023) for GAAH model which hosts a single particle mobility edge in the non-interacting limit over a range of parameters. The mobility edge persists in the SP excitations even in the presence of interaction in the nonergodic extended (NEE) phase that exists over a broad range of many-body energy in the finite-sized systems accessed by ED. The NEE phase in the GAAH model emerges as a result of the interaction-induced mixing of delocalized and localized single-particle excitations separated by a mobility edge in the noninteracting systems Modak and Mukerjee (2015); Ghosh et al. (2020); Ganeshan et al. (2015); Deng et al. (2017); Roy et al. (2023). However, in the EAAH model, the MBC phase is obtained due to interaction-induced mixing of the single-particle critical states, in the absence of any SP mobility edge in the noninteracting limit. Thus the feature of the non-interacting critical phase, namely the delocalized nature of all SP excitations, tends to persist even in the MBC phase. Moreover, the MBL phase in the GAAH model has been argued Modak and Mukerjee (2015); Ghosh et al. (2020); Roy et al. (2023) to be an outcome of the mechanism called as the “MBL proximity effect” Nandkishore (2015). On the contrary, the MBC phase in the EAAH model arises due to the stability of the critical phase to interaction, at least in the finite systems accessed by ED.
IV Dynamical transitions and Widom lines in Fock-space self energy and inverse participation ratio
[…]
The FS sites are arranged in slices such that any site in a particular slice is connected to the sites of the nearest slices by a single FS hopping, which is shown in Fig. 4. The local structure of the FS lattice allows one to implement a recursive Green’s function method MacKinnon (1980); MacKinnon and Kramer (1983); Lee and Fisher (1981); Vergés (1999), which has been recently applied to the FS lattice of similar models with random Sutradhar et al. (2022) and quasiperiodic disorders Roy et al. (2023). We compute the elements of the retarded FS propagator $G(E)=[E+i\eta-\tilde{H}]^{-1}$ at the middle of the energy spectrum $E=0$, corresponding to the infinite-temperature state. Here we choose a broadening parameter $\eta(\mu,\lambda)=\sqrt{2\pi}\sigma_{E}/(\sqrt{L}\mathcal{N}_{F})$ Welsh and Logan (2018); Logan and Welsh (2019), i.e., the mean energy-level spacing, which is a function of parameters $\mu$ and $\lambda$.
[…]
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 5: Fock-space self-energy for noninteracting systems: Typical values of self-energy $\Delta_{t}$ across (a) the delocalization-localization transition as a function of $\lambda$ for $\mu=0.5$, (b) delocalization-critical transition as a function of $\mu$ for $\lambda=1.0$, (c) critical-localization transitions as a function of $\lambda$ for $\mu=1.5$, and (d) localization-critical transition as a function of $\mu$ for $\lambda=3.5$. The phase transition points are denoted by the vertical solid lines. In figures (c) and (d), the vertical dashed lines denote the peak/dip in $\Delta_{t}$ across the Widom-like lines in Fig.1(a) inside the critical and localized phases, respectively.
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 6: Fock-space IPRIPR for noninteracting systems: IPR across (a) the delocalization-localization transition as a function of $\lambda$ for $\mu=0.5$, (b) delocalization-critical transition as a function of $\mu$ for $\lambda=1.0$, (c) critical-localization transitions as a function of $\lambda$ for $\mu=1.5$, and (d) localization-critical transition as a function of $\mu$ for $\lambda=3.5$. Vertical solid lines denote the points of phase transitions. In figures (c) and (d), the vertical dashed lines denote the peak/dip in IPR across the Widom-like lines in Fig.1(a) inside the critical and localized phases, respectively.
IV.1 Noninteracting systems
We begin with the discussion of $\Delta_{t}$ in the noninteracting system. In Fig. 5(a-d), we show the variation of $\Delta_{t}$ across the phase transitions along four chosen cuts on the phase diagram of Fig. 1(a) to study $\Delta_{t}$ across delocalized to localized, delocalized to critical, and critical to localized phase transitions. In Fig. 5(a), $\Delta_{t}$ is shown across the delocalization-localization phase transition (at $\lambda=2$). As evident in Figs. 5(a) and (b), $\Delta_{t}\sim\mathcal{O}(1)$ and does not vary much with system size $L$ in the delocalized phase. On the contrary, $\Delta_{t}$ rapidly decreases with $L$ in the localized phase. The delocalization-critical phase transition at $\mu=1$ for $\lambda=1$ is shown in 5(b). $\Delta_{t}$ is non-monotonic in the critical phase, exhibiting a sudden drop near the transition. In the critical phase, $\Delta_{t}$ shows an anomalous non-monotonic dependence, and thus no systematic scaling with $\mathcal{N}_{F}$, at least up to the largest system sizes accessed here ($L=20$). Fig. 5(c) and Fig. 5(d) show variation of $\Delta_{t}$ across the critical-localization and localization-critical phase transitions as a function of $\lambda$ ($\mu=1.5$), and as a function of $\mu$ ($\lambda=3.5$), respectively. $\Delta_{t}$ in the critical phase again shows a non-monotonic system size and/or parameter dependence.
Remarkably, there is a distinct dip, indicating stronger localizing tendency, at $\mu=1$ in the localized phase in Fig. 5(d). We have verified that the dip is present along the $\mu=1$ line as a function of $\lambda$ throughout the localized phase. The $\mu=1$ line, emanating from the delocalized-localized-critical triple point, is a continuation of the dip in $\Delta_{t}$ [Fig. 5(d)] near the delocalized-critical phase boundary, as elucidated in Fig.1(a). This marked non-monotonic feature also appears as a peak in IPR as we discuss below. The locus of distinct dips (peaks) in $\Delta_{t}$ (IPR) along the $\mu=1$ line also persists in the MBL phase in the interacting EAAH model, as shown in Fig.1(b), and discussed later. Generally, such a line of dips or peaks in physical quantities within a phase as an extension of a phase transition line or critical point is reminiscent of the well-known Widom or Fisher-Widom lines seen in supercritical fluid Xu et al. (2005); Franzese and Stanley (2007); Simeoni et al. (2010); Luo et al. (2014); Sordi and Tremblay (2024) or near the metal-insulator critical point Sordi et al. (2012, 2013); Vučičević et al. (2013) in the equilibrium thermodynamic phase diagram. There, the Widom line, manifests as a peak in the correlation length or peaks/dips in various thermodynamic quantities like the specific heat, compressibility etc., may indicate a crossover or weak singularity or even in the presence of some hidden critical point Xu et al. (2005).
Curiously, our work reveals the presence of Widom-like lines even in the dynamical phase diagram, presumably indicating a crossover within the localized phase due to the antecedent delocalized-critical phase transition line along $\mu=1$. Moreover, a weaker non-monotonic feature, in the form of a dip or peak in $\Delta_{t}$ [Fig. 5(c)] and dip in IPR [Fig.6(c)], is also seen as a function of $\lambda$ within the critical phase along $\lambda=2$ line, as shown in Fig.1(a). The $\lambda=2$ line is in the critical phase is an extension the delocalized-localized phase boundary as well as the self-dual line. This Widom-like line, shown in Fig.1(b), inside the critical phase becomes even more prominent in the interacting case, as we discuss in the next section.
[…]
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)           Refer to caption(e)           Refer to caption(f)
Figure 7: Finite-size dependence of $IPR$ for noninteracting systems: (a) Log-log plots of $IPR$, after linear fitting, as a function of FS dimension $\mathcal{N}_{F}$ across the delocalization-localization transition as a function of $\lambda$ for $\mu=0.5$. (b) Spectral dimension $D_{2}$ extracted from $IPR=A_{I}\mathcal{N}_{F}^{-D_{2}}$ across the delocalized-localized transition in (a). (c,d) and (e,f) Similar figures as (a,b) but for the critical-localized transition as a function of $\lambda$ for $\mu=1.5$ and the delocalized-critical transition as a function of $\mu$ for $\lambda=1.0$, respectively. Blue squares in figures (b,d,f) indicate $D_{2}$ obtained from the finite-size scaling analysis (not shown here) in the delocalized and localized phases. In figure (d) the vertical dashed line corresponds to the Widom line along $\lambda=2$ in the critical phase. Vertical solid lines denote phase transitions in all the plots.
We extract the coefficient $A_{I}$ and the fractal dimension $D_{2}$ from the $IPR\simeq A_{I}\mathcal{N}_{F}^{-D_{2}}$ through a linear fitting of the data shown in Fig. 7(a,c,e). From this analysis we conclude that $D_{2}\approx 1$, $A_{I}<1$ in the delocalized phase, whereas $0<D_{2}<1$, $A_{I}>1$ in the localized phase. The critical phase exhibits IPR in between the delocalized and localized phases, namely $0<D_{2}<1$ like the localized phase, but $A_{I}>1$ as in the delocalized phase. We also extract similar coefficients and exponents, $A_{s}$ and $D_{s}$, respectively, from $\Delta_{t}$, as summarized in Table 1 (see also Appendix B). However, $\Delta_{t}$ is rather irregular as a function of $\mathcal{N}_{F}$ in the critical phase, unlike the IPR [Fig. 7(c)] and, as a result, we could not extract $A_{s}$ and $D_{s}$ reliably in the critical phase.
In Figs.7(b,d,f), we also plot $D_{2}$ obtained from the finite-size scaling analysis of the IPR in the delocalized and localized phases approaching the delocalized-localized, localized-critical and delocalized-critical transitions. The finite-size scaling ansatz Macé et al. (2019); Garcia-Mata et al. (2017); Roy and Logan (2021); Sutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025), discussed in detail in the next section for the interacting EAAH model, assumes a volumic scaling for the delocalized phase and a linear scaling for the localized phase. As evident, $D_{2}$ estimated from finite-size scaling in the localized phase [Figs.7(b,d)] agrees reasonably with the $D_{2}$ extracted by directly the fitting the data for $IPR$ as a function of $\mathcal{N}_{F}$ in the localized phase in Figs.7(a,c). The volumic finite-size scaling in the delocalized phase assumes $D_{2}=1$ by default. This value is only expected in the asymptotic thermodynamic limit $\mathcal{N}_{F}\to\infty$. In Figs.7(b,f), we see that $D_{2}$ extracted directly from the data in Figs.7(a,e) for the delocalized phase deviates by $\sim 10-20\%$ from the asymptotic $D_{2}=1$ even deep inside the delocalized phase for the system sizes accessed in this work. Due to the non-monotonic, and somewhat irregular, behavior of the $IPR$ as a function of $\mathcal{N}_{F}$ and the parameters $\lambda$ or $\mu$ in the critical phase, $D_{2}$ could not be extracted from finite-size scaling in the critical phase, using either volumic or linear scalings.
Note that within the sizes that we are able to study $(L=28)$, the many-body noninteracting localized phase is multifractal in the FS, unlike in the case of Anderson localization in real-space which has a fractal dimension zero. Interestingly, we find that the change of $A_{I}$ (not shown) and $D_{2}$ [Figs.7(d,f)] across the critical-localized and the delocalized-critical phase transitions are much more abrupt than those across the delocalized-localized transition, e.g., for $D_{2}$ in Fig.7(b). The abrupt jumps of values of $D_{2}$ in Figs.7(d,f) coincide very well with the critical-localized and delocalized-critical transitions estimated from other diagnostics Wang et al. (2021). This provides us with confidence to study these quantities further the interacting systems, as discussed in the next section.
IV.2 Interacting systems
Here we discuss our results for FS self-energy and IPR in the interacting EAAH model with $V=1$. In Fig. 8(a-d), we show the variation of $\Delta_{t}$ of interacting systems for the same choices of parameters as in Fig. 5(a-d), respectively, involving all three phases [see Fig. 1(b)]. In Fig. 8(a), $\Delta_{t}$ is shown across the ergodic-MBL phase transition, where $\Delta_{t}\sim\mathcal{O}(1)$ in the ergodic phase and rapidly decreases with $L$ in the MBL phase. The ergodic-MBC phase transition at $\mu=1$ is shown in 8(b), where $\Delta_{t}$ abruptly drops near the transition, like in the non-interacting system, and then increases again deeper in the MBC phase. Fig. 8(c) and Fig. 8(d) show variation of $\Delta_{t}$ across the MBC-MBL and MBL-MBC phase transitions, respectively. $\Delta_{t}$ in the MBC phase shows an anomalous $L$-dependence. In Fig. 8(c), we see an enhanced delocalization tendency around $\lambda=2$ as a broad peak in the MBC phase. The peak in $\Delta_{t}$ appears as a dip in IPR, as we discuss below. The peak is indicative of the Widom line in the MBC phase, as shown in Fig.1(b). However, unlike in the non-interacting model where the Widom line comes out from the triple point $\lambda=2,\mu=1$, the locus of the peak in the MBC phase emanates from the ergodic-MBC phase boundary around $\lambda\approx 1.5,\mu=1$, away from the triple point ($\lambda\approx 3,\mu=1$) in the interacting EAAH model. However, deeper inside the MBC phase, the Widom line eventually merges with the $\lambda=2$ line, which is the Widom line in the critical phase of the non-interacting EAAH. Similarly, in the MBL phase, $\Delta_{t}$ has a minimum along the $\mu=1$ Widom line, implying stronger localization, as shown in Fig. 8(d). A recent numerical work, based on the supervised machine learning approach to detect various phases of our many-body quasi-periodic system Ahmed and Roy (2025), has also indicated the possibility of the existence of crossovers around the lines $\lambda=2$ and $\mu=1$ within the MBC and MBL phases, respectively .
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 8: Fock-space self-energy for interacting systems: Typical values of self-energy $\Delta_{t}$ across (a) the ergodic-MBL transition as a function of $\lambda$ for $\mu=0.5$, (b) ergodic-MBC transition as a function of $\mu$ for $\lambda=1.0$, (c) MBC-MBL transitions as a function of $\lambda$ for $\mu=1.5$, and (d) MBL-MBC transition as a function of $\mu$ for $\lambda=3.5$, for increasing system sizes $L$ . In figures (c) and (d), the vertical dashed lines denote the non-monotonic dependence of $\Delta_{t}$ and IPR on the parameters in MBC and MBL phases, respectively. The solid vertical lines denote the points of the phase transition.
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 9: Fock-space IPRIPR for interacting systems: IPRIPR across (a) the ergodic-MBL transition as a function of $\lambda$ for $\mu=0.5$, (b) ergodic-MBC transition as a function of $\mu$ for $\lambda=1.0$, (c) MBC-MBL transitions as a function of $\lambda$ for $\mu=1.5$, and (d) MBL-MBC transition as a function of $\mu$ for $\lambda=3.5$, for increasing system sizes $L$. In figures (c) and (d), the vertical dashed lines denote the peak/dip of $\Delta_{t}$ across the Widom lines in Fig.1(b). The solid vertical lines denote the phase transitions.
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)           Refer to caption(e)           Refer to caption(f)
Figure 10: Finite-size dependence of $IPR$ for interacting systems: (a) Log-log plots of $IPR$, after linear fitting, as a function of FS dimension $\mathcal{N}_{F}$ across the ergodic-MBL transition with $\lambda$ at $\mu=0.5$. (b) Spectral dimension $D_{2}$ extracted from $IPR=A_{I}\mathcal{N}_{F}^{-D_{2}}$ for the ergodic-MBL transition in (a). (c,d) and (e,f) Similar figures as (a,c) for MBC-MBL transition as a function of $\lambda$ at $\mu=1.5$ and ergodic-MBC transition as function of $\mu$ at $\lambda=1.0$. The vertical solid lines denote the phase transitions in all plots. $D_{2}$ obtained from the finite-size scaling analysis of Fig. 11 is shown with blue squares in figures (b,d,f). In figure (d) the vertical dashed line corresponds to the Widom line along $\lambda=2$ in the critical phase.
We now discuss our results for FS IPR, which is shown in Fig. 9(a-d) for the same transitions shown in Fig. 8(a-d), respectively. Here we employ exact diagonalization of the interacting Hamiltonian. As a result, we are able to study the system only up to a system size $L=16$. In all the three phases, i.e., ergodic, MBL and MBC, the $IPR$ decreases with increasing $L$. The non-monotonicity across the Widom lines [Fig.1(b)] are again visible in IPR in Figs.9(c) and (d). The $IPR$ as a function of FS dimension $\mathcal{N}_{F}$ is shown in Fig. 10(a,c,e) for the same parameters as in Fig. 9(a,b,c), respectively. The coefficient $A_{I}$ and $D_{2}$ [Figs.9(b,d,f)] are obtained through a linear fitting of the data in Fig. 10(a,c,e). We find that $(D_{2}\approx 1,A_{I}>1)$, $(0<D_{2}<1,A_{I}<1)$ and $(0<D_{2}<1,A_{I}>1)$ in the ergodic, MBL and MBC phases, respectively. Similar results are obtained for $A_{s}$ and $D_{s}$ extracted from $\Delta_{t}$ (Table 1, see also Appendix B), except in the MBC phase where we could not extract $A_{s}$ and $D_{s}$ to the non-monotonic dependence of $\Delta_{t}$ on $\mathcal{N}_{F}$. The signature of the Widom line along $\lambda=2$ is visible as a broad peak in $D_{2}$ as a function of $\lambda$ in Fig. 10(d). The solid vertical lines in Fig. 10(b,d,f) denote the phase transitions at the critical values $\lambda_{c}\approx 3.2$ ($\mu=0.5$), $\lambda_{c}\approx 3.2$ ($\mu=1.5$) and $\mu_{c}\approx 1.0$ ($\lambda=1$) obtained from the finite-size scaling analysis, as we discuss next.
For the finite-size scaling analysis of $IPR$ across the different transitions of the interacting system, we consider the volumic and linear scaling ansatzs Macé et al. (2019); Garcia-Mata et al. (2017); Roy and Logan (2021) which are given by,
$-\ln\frac{I_{t}}{I_{c}}=\begin{cases}\mathcal{F}_{vol}\big(\frac{\mathcal{N}_{F}}{\Lambda}\big)&~~:~\text{Ergodic}\\ \mathcal{F}_{lin}\big(\frac{\ln\mathcal{N}_{F}}{\xi}\big)&~~:~\text{MBC, MBL}.\end{cases}$ (6)
Figure 11: Scaling collapse of $IPR$ for interacting systems: (a,b) Ergodic-MBL transition: Volumic scaling and linear scaling of $IPR$ in the ergodic and MBL phases, respectively, for the ergodic-MBL transition ($\lambda_{c}=3.2$) as a function of $\lambda$ for $\mu=0.5$ corresponding to Fig. 10(a,b). (c,d) MBL-MBC transition: Linear scaling of $IPR$ in the MBC and MBL phases, respectively for the MBC-MBL transition $(\lambda_{c}=3.2)$ as a function of $\lambda$ for $\mu=1.5$, corresponding to Fig. 10(c,d). (e) Ergodic-MBC transition: Volumic scaling of $IPR$ in ergodic phase for the ergodic-MBC phase transition $(\mu_{c}=1.0)$ as a function of $\mu$ for $\lambda=1.0$, corresponding to Fig. 10(e,f). Insets show the KT-like divergence of the non-ergodic volume $\Lambda$ and the power-law divergence of correlation length $\xi$ near phase transitions in case of volumic and linear scaling, respectively.
Here $I_{t}$ is the typical value of IPR that depends on $\mathcal{N}_{F}$ and other parameters like $\lambda,~\mu$; $I_{t}=I_{c}$ at the critical point. In the thermodynamic limit $\mathcal{N}_{F}\to\infty$, we expect $I_{t}\sim\mathcal{N}_{F}^{-D_{2}}$ ($I_{c}\sim\mathcal{N}_{F}^{-D_{2c}}$) with an exponent $D_{2}$ ($D_{2c}$). According to the scaling ansatz, the scaling form inside the ergodic phase is volumic controlled by a ‘nonergodic volume’ $\Lambda$ that diverges when the ergodic system approaches a phase transition towards a nonergodic phase. On the other hand, in the nonergodic phases, like MBC and MBL phases, we assume a linear scaling form controlled by a length scale $\xi$ that diverges as the system approaches a non-ergodic to ergodic phase transition. The scaling ansatzs of Eq.(6) can also be used to study the FS IPR in the non-interacting EAAH model across the delocalized-localized, critical-localized, and delocalized-critical transitions. For example, we have used the volumic scaling in the delocalized phase and the linear scaling in the localized and critical phases to extract the fractal dimension $D_{2}$ (see the discussion below), as shown in Figs.7(b,d,f).
The asymptotic behavior Roy and Logan (2021); Sutradhar et al. (2022) of volumic scaling function for $IPR$ [Eq.(6)] in the ergodic phase for $\mathcal{N}_{F}\gg\Lambda$ is given by $\mathcal{F}_{vol}(\mathcal{N}_{F}/\Lambda)\sim(1-D_{2c})\ln{(\mathcal{N}_{F}/\Lambda)}$ where we assume $D_{2}=1$ for the ergodic phase. The asymptotic form of the linear scaling function in the limit $\ln{N_{F}}\gg\xi$ in both the nonergodic phases is given by $\mathcal{F}_{lin}(\ln\mathcal{N}_{F}/\xi)\sim-D_{2c}\ln{(\mathcal{N}_{F})}/\xi$ where $\xi=D_{2c}/(D_{2c}-D_{2})$. Thus, the values of $\xi$ extracted as a function of $\lambda$ or $\mu$ from the finite-size scaling collapse in the MBL and MBC phases, can be used to obtain the fractal dimension $D_{2}$ in these non-ergodic phases. We also use the above asymptotic scaling functions to obtain estimates of $D_{2c}$ by approaching the transitions from either sides. For example, for the ergodic-MBL transition we extract $D_{2c}$ using the volumic scaling from the ergodic side, and using the linear scaling from the MBL side. Similarly, we use two different linear scaling collapses for the MBC-MBL transition, one from the MBC, and the other from the MBL side. In all the cases, where reliable scaling collapses can be obtained and $D_{2c}$ can be estimated independently in two different sides of the transition, $D_{2c}$ values match. This gives confidence in our finite-size scaling procedure.
Ergodic-MBL transition.– In Figs. 11(a,b) we show the volumic and linear finite-size scaling collapses in the ergodic and MBL phases, respectively, for the ergodic-MBL transition as a function of $\lambda$ at $\mu=0.5$, as shown in Fig. 10(a,b). The quality of the collapse is estimated from the $\chi^{2}$ difference between the data points and fitted scaling functions. The best quality of the volumic scaling collapse shown in Fig. 11(a) is obtained for the critical value $\lambda_{c}=3.2$. From the extracted scaling function in the ergodic phase [Fig. 11(a)], we find $\mathcal{F}_{vol}(\mathcal{N}_{F}/\Lambda\gg 1)\sim 0.83\ln{(\mathcal{N}_{F}/\Lambda)}$, implying $D_{2c}\approx 0.17$. The extracted non-ergodic volume follows $\Lambda\sim\exp{[b(\lambda_{c}-\lambda)^{-0.4}]}$ [Fig. 11(a)(inset)] with $b\sim\mathcal{O}(1)$. In the MBL phase, from Fig. 11(b), we find $\mathcal{F}_{lin}(\ln\mathcal{N}_{F}/\Lambda)\sim-0.17\ln{(\mathcal{N}_{F})}/\xi$, i.e., with the same value of $D_{2c}\simeq 0.17$ as obtained from the ergodic side. The extracted length scale $\xi$ [Fig. 11(b)(inset)] diverges as $\xi\sim|\lambda-\lambda_{c}|^{-0.5}$. Similar exponents have been reported in a recent study on ergodic-MBL transition in quasiperiodic AAH chain using imaginary part of FS self-energy Ghosh et al. (2025).
MBC-MBL transition.– Figs. 11(c,d) demonstrate the finite-size scaling analysis for the MBC-MBL phase transition of Fig. 10(c,d) as a function of $\lambda$ at $\mu=1.5$. We find that the linear scalings are governed by correlation lengths $\xi_{1}$ and $\xi_{2}$ in the MBC and MBL phases, respectively. The analysis leads to $\lambda_{c}=3.2$, $D_{2c}\approx 0.3$ and $\xi_{1},\xi_{2}\sim|\lambda-\lambda_{c}|^{-0.4}$.
Ergodic-MBC transition.–Fig. 11(e) shows the volumic scaling in the ergodic phase corresponding to ergodic-MBC transition with $\mu$ at $\lambda=1.0$, as depicted in Fig. 10(e,f). The data collapse in the ergodic phase indicates that $\mathcal{F}_{vol}(\mathcal{N}_{F}/\Lambda)\sim 0.44\ln{(\mathcal{N}_{F}/\Lambda)}$ with $\mu_{c}=1$, $D_{2c}\approx 0.56$ and $\Lambda\sim\exp{[b_{2}(\mu_{c}-\mu)^{-0.9}]}$ with $b_{2}\sim\mathcal{O}(1)$. Based on the $\chi^{2}$ analysis, a reasonable collapse of the data in the MBC phase was not possible to achieve. Our scaling analysis of $IPR$ shows that the asymptotic form of the volumic scaling, governed by $\Lambda$, in the ergodic phase can change substantially depending on whether the ergodic-to-non-ergodic transition is approached from ergodic to a MBL phase or a MBC phase.
Scaling ansatz Sutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025) similar to Eq.(6) can be written for the FS self-energy $\Delta_{t}$. Using these scaling ansatz for $\Delta_{t}$, we have also performed the finite-size scaling analysis for the data of self-energy across ergodic-MBL, ergodic-MBC and MBC-MBL transitions in Figs.8(a,b,c,d). The analysis across the ergodic-MBL transition, as discussed in Appendix B, leads to qualitatively similar results as in the earlier studies on random and other quasiperiodic systemsSutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025). However, due to more irregularity and non-monotonicity of $\Delta_{t}(\mathcal{N}_{F},\lambda,\mu)$ compared to that in $IPR$ in the MBC phase, we could not obtain reliable data collapse across the transitions to MBC phase. Hence, we have only discussed the finite-size scaling analysis of $IPR$ above. The distributions of IPR and $\Delta_{I}$ for different phases show distinct features which are discussed in Appendix C.
V Fock-space localization length
In this section, we analyze the off-diagonal elements $G_{IJ}(E=0)$ of the FS propagator Roy et al. (2023); Ghosh et al. (2025) in the ergodic, MBL and MBC phases of the interacting EAAH model, and compare them with the phases of the non-interacting model. We calculate the geometric mean or the typical value, defined as $G(r_{IJ})=\text{exp}[\langle\ln G_{IJ}\rangle]$, where $\langle$$\rangle$ denotes an average over $\phi$ and all the off-diagonal elements $G_{IJ}$ for a pair of FS sites are at a hopping distance $r_{IJ}$. The latter is the minimum number of nearest-neighbor hops to go from $I$ to $J$ in the middle slice. The number of realizations used for $\phi$ are $2000,1000,400,200,100$ for $L=12,14,16,18,20$, respectively. Plots of $\ln[G(r_{IJ})]$ as a function of $r_{IJ}$ are shown in Figs. 12(a-c) and Figs. 12(d-f) for the phases of the interacting and non-interacting EAAH model. In all the phases, the plots show a linear regime with negative slope, i.e., $\ln[G(r_{IJ})]\propto-r_{IJ}$, before it deviates from linearity depending on $L$. This deviation from the linear regime corresponds to rare hopping distances and associated multiple length scales in FS Roy et al. (2023); Ghosh et al. (2025). The linear regime implies the existence of a decay length $\xi_{F}(L)$, dubbed as the ‘FS localization length’ Roy et al. (2023); Ghosh et al. (2025), such that $G(r_{IJ})\sim\text{exp}[-r_{IJ}/\xi_{F}(L)]$.
     Refer to caption(a)           Refer to caption(d)           Refer to caption(b)           Refer to caption(e)           Refer to caption(c)           Refer to caption(f)
Figure 12: Decay of two-point correlations in Fock-space: (a-c) Decay of typical values $G(r_{IJ})$ of correlations with FS-hopping distance $r_{IJ}$ in the ergodic ($\lambda=1.0,~\mu=0.5$), MBL ($\lambda=5.0,~\mu=0.5$) and MBC ($\lambda=1.0,~\mu=1.5$) phases, respectively, of the interacting EAAH model. (d-f) Similar plots in the delocalized, localized and critical phases, respectively, of noninteracting model.
In the ergodic phase, $\xi_{F}(L)$ rapidly increases with $L$ [Fig. 12(a)]. In comparison, $\xi_{F}(L)$ increases much more slowly with $L$ in the non-interacting delocalized phase [Fig. 12(d)]. Similarly, in the MBC phase, $\xi_{F}(L)$ shows more rapid increases with $L$ [Fig. 12(c))] compared to $\xi_{F}(L)$ in the non-interacting critical phase [Fig. 12(f)]. The faster increase of the FS localization length with $L$ in the interacting model presumably reflects the effect of interaction that induces more correlation in the FS lattice, leading to a greater tendency to localize. In contrast to the ergodic (delocalized) and MBC (critical) phases, in the MBL (localized) phase, $\xi_{F}(L)$ for different values of $L$ overlap up to a certain $L$-dependent value of $r_{IJ}$ and then show a weak dependence on $L$ [Fig. 12(b) (Fig. 12(e))]. The non-interacting AL phase shows a more stable linear regime and weaker dependence on $L$.
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 13: Fock-space localization length: (a-b) FS localization length $\xi_{F}$ extracted from the decay of $G_{t}(r_{IJ})$ across the ergodic-MBL transition as a function of $\lambda$ for $\mu=0.5$, and ergodic-MBC transition as function of $\mu$ for $\lambda=1.0$ for the interacting EAAH model. (c-d) $\xi_{F}$ across the delocalization-localization transition as a function of $\lambda$ for $(\mu=0.5)$ and delocalization-critical transitio as a fucntion of $\mu$ for $\lambda=1.0$ for the non-interacting EAAH model. Vertical dashed lines denote the transitions in the non-interacting and interacting systems.
Plots in Figs. 13(a,b) show the variation of $\xi_{F}$ with $\lambda$ or $\mu$ for increasing $L$ across the ergodic-MBL and ergodic-MBC transitions, respectively. The FS localization length $\xi_{F}$ is almost independent of $L$ in the MBL phase, whereas it depends strongly on $L$ in the ergodic phase, and weakly in the MBC phase. Similarly, plots for the non-interacting system are shown in Figs. 13(c,d) across the delocalization-localization and delocaliazation-critical phase transitions, respectively. Deep in the localized phase, $\xi_{F}$ is independent of $L$, whereas it increases with $L$ in the delocalized and critical phases, albeit slowly compared to $\xi_{F}(L)$ in the ergodic and MBC phases of the interacting EAAH model. The non-monotonic localization properties across the ergodic (delocalized) to MBC (critical) phase transition are nicely captured by $\xi_{F}$, similar to $\Delta_{t}$ and IPR [Figs.8,9].
VI Widom-like line in the presence of random potential
In this short section, we make an attempt to find Widom-like line in presence of random disorder. It is not clear at the moment how to find a Widom-like line in a truly random system as appearance of Widom line would typically need presence of a special critical line or a triple point. However, one can think of a single particle model with both a regular hopping $(J=1)$ and quasiperidoc hopping (of strength $\mu$), like in EAAH model plus a random onsite potential (of strength $W$) instead of the quasiperiodic potential. In that case, for $W=0$ there will be a transition from delocalized to critical at $\mu=1$, according to Fig. 1(a) in the main text. For finite non-zero $W$ all single particle states become localized for all values of $\mu$. Interestingly, we find that the states show the strongest localization along the $\mu=1$ line for $W>0$. Hence, one can say that a Widom-like line emanates from $(\mu=1,W=0)$ point and continues for finite $W$ (see Fig. 14). Presumably this $\mu=1$ line will continue to persist even in the stipulated MBL phase in presence of interaction.
     Refer to caption(a)           Refer to caption(b)
Figure 14: Tight-binding model with quasiperiodic hopping and random onsite potential: (a) For $W\neq 0$, $IPR$ as a function of $\mu$ shows a peak at around $\mu=1$ tracing the presence of Widom-like line. (b) Phase diagram: In absence of onsite random potential $(W=0)$, the single particle system undergoes delocalized-to-critical phase transition at $\mu=1$ denoted by the black dot. In presence of random potential $(W\neq 0)$, all the single particle eigenstates become localized. However, a Widom-like line denoted by the red dot-dashed line emanates from the critical point $\mu=1$ showing a line of the strongest localization as demonstrated in figure (a).
VII Discussion and Conclusions
To summarize, we provide a characterization of the MBC (critical) phase in the interacting (non-interacting) extended Aubry-André-Harper (EAAH) model through localization properties of excitations in real and Fock space, and eigenstate IPR. We also analyze critical properties of the transitions between the MBC (critical) phase and other phases, like ergodic (delocalized) and MBL (localized). We show that the MBC (critical) phase is described by multifractal scaling of FS IPR, as well as a linear finite-size scaling ansatz for IPR with a diverging length scale near the MBC (critical) to ergodic (delocalized) and MBL (localized) transition. However, the MBC (critical) phase hosts delocalized single-particle excitations in real-space and exhibits a system-size dependent FS localization length, like the ergodic (delocalized) phase. Most interestingly, we show that the ergodic (delocalized)- MBC (critical) phase transition line, even after terminating at triple point, continues as a Widom-like line in the MBL (localized) phase. Similarly, another Widom line, emanating from the ergodic-MBC phase boundary (triple point) persists in the MBC (critical) phase. These Widom lines are manifested in the FS localization properties, namely as peaks/dips (dips/peaks) in the FS self energy (FS IPR).
In the thermodynamic limit, the NEE phases, like the MBC phase, is expected to be more fragile than the MBL phase. The stability of the MBL phase in the thermodynamic limit for systems with random disorder has been called to question due to the possibility of long-range resonances and avalanche instabilities from rare weak-disorder regions in larger system sizes, not accessible via exact diagonalization $(L\leq 24)$ Sels and Polkovnikov (2021); Sels (2022); Crowley and Chandran (2022); Long et al. (2023); Morningstar et al. (2022); Sierant and Zakrzewski (2022). However, unlike the systems with random disorder, the quasiperiodic systems are not susceptible to the usual avalanche instability De Roeck and Huveneers (2017) due to the absence of rare weak-disorder regions Agarwal et al. (2017). Hence the MBC phases, like the MBL phase, in the quasiperiodic EAAH model could be more robust. The mechanism of obtaining NEE states from the single-particle critical states is a new one and may produce more stable NEE states than found in previous works Li et al. (2015); Deng et al. (2017) on the quasiperiodic GAAH model. Nevertheless, the issue with extrapolating the existence of MBC phase, based on the small finite-size numerics, to the thermodynamic limit is expected to be more complex than that for the MBL phase. In this regard, it might be useful to apply matrix-product state (MPS) or density-matrix renormalization group based approaches for the EAAH model to access larger system sizes, as has been attempted Pomata et al. (2023) for the GAAH model. However, the MBC states have volume-law entanglement and thus harder to describe through MPS-based approaches than the MBL phase.
The EAAH model and a related model have been realized experimentally with superconducting circuits Li et al. (2023) and ultra-cold atoms Xiao et al. (2021), respectively. In future, with development and improvement of energy-resolved spectroscopies Kollath et al. (2007); Stewart et al. (2008); Perali et al. (2011); Jiang et al. (2011) for the cold-atomic systems, it might be possible to probe LDOS that would be able to track the localization phenomena in real-space. It will also be interesting to look at the short-range and long-range resonances Crowley and Chandran (2022); Long et al. (2023); Morningstar et al. (2022); Garratt and Roy (2022); Garratt et al. (2021) in the MBC phase, as well as entanglement growth and the related four-point correlations Roy (2022) in Fock-space. A better understanding of the nature of the Widom lines and the mechanism behind their origin in the dynamical phase diagram of the EAAH model, and realizations of such lines in other models would be interesting research directions to pursue in future.
Acknowledgements
SB acknowledges support from CRG, SERB (ANRF), DST, India (File No. CRG/2022/001062) and STARS, MoE, Govt. of India (File. No. MoE-STARS/STARS-2/2023-0716)
Appendix A Energy level-spacings and spectral form factor
We discuss here the analysis of distributions $P(s)$ of consecutive many-body energy level-spacings $s$, normalized by the (arithmetic) mean level spacing, in the ergodic, MBL and MBC phases of the interacting EAAH model. The nearest neighbor level-spacing distributions for ergodic, MBL and MBC phases are shown in Fig. 15(a-c), respectively, for increasing $L$. For the largest system size $L=16$, we fit the data to the Brody distribution given by $P(s)=As^{a}\exp(-As^{a+1}/(a+1))$ where $A=(a+1)\Gamma(\frac{a+2}{a+1})^{(a+1)}$. $a=1$ and $a=0$ correspond to GOE and Poisson distributions, respectively. We find $a\approx 0.94$, $0.02$ and $0.26$ for the choices of parameters in Fig. 15(a-c), respectively. The plots indicate that $P(s)$ approaches GOE distribution in the ergodic phase and Poisson distribution in the MBL phase Atas et al. (2013). Moreover, for the MBC phase $0.15<a<0.26$, implying distributions intermediate between GOE and Poisson, where maximum value of $a\approx 0.26$ appears along the Widom line ($\lambda\approx 2$) in the MBC phase.
     Refer to caption(a)           Refer to caption(d)           Refer to caption(b)           Refer to caption(e)           Refer to caption(c)           Refer to caption(f)
Figure 15: Energy level statistics and correlations in the interacting EAAH model: (a-c) Level spacing distributions $P(s)$ in the ergodic $(\mu=0.5,\lambda=1.0)$, MBL $(\mu=0.5,\lambda=5.0)$, MBC $(\mu=1.5,\lambda=2.0)$ phases, respectively. The dashed lines are fit to the Brody distribution to the data for $L=16$ . (d-f) Spectral form factor (SFFSFF) of the interacting systems for the same parameters as in figures (a-c), respectyively. For calculation of $SFF$, the number of mid-spectrum states $N_{st}=1000$.
$P(s)$ captures the short-range correlations of the many-body energy spectrum. There is another quantity, namely the spectral form factor, that includes all possible correlations in the spectrum. One first defines a function $Z(\tau)=\sum_{n=1}^{N_{st}}e^{-iE_{n}\tau}$ of fictitious time $\tau$ including $N_{st}$ number of eigenstates. Then the spectral form factor can be written as Haake (1991)
$SFF(\tau)$ $=$ $\langle Z^{*}(\tau)Z(\tau)\rangle,$ (7)
$=$ $N_{st}+\langle\sum\limits_{m\neq n}e^{-i(E_{m}-E_{n})\tau}\rangle,$
where $\langle$$\rangle$ stands for an average over some ensemble, e.g., realizations of $\phi$ in the EAAH model. At $\tau=0$, $SFF=N_{st}^{2}$. For very small $\tau\ll\tau_{D}$ compared to the Thouless time $\tau_{D}$, $\langle Z^{*}(\tau)Z(\tau)\rangle=\langle Z^{*}(\tau)\rangle\langle Z(\tau)\rangle$ due to absence of spectral correlation and $SFF$ decreases showing non-universal model-specific spectral features. At $\tau=\tau_{H}$, the Heisenberg time, $\tau$ becomes comparable to inverse of mean level spacing. As a result, the second term in Eq. 7 vanishes on average and a plateau at $SFF=N_{st}$ appears. For $\tau_{D}<\tau<\tau_{H}$, $SFF$ shows a ramp indicating the development of spectral correlation such that $\langle Z^{*}(\tau)Z(\tau)\rangle\neq\langle Z^{*}(\tau)\rangle\langle Z(\tau)\rangle$. This region shows universal features and does not appear in the absence of any spectral correlation. In Fig. 15(d-f) rescaled $SFF/N_{st}$ is plotted for the three phases corresponding to Fig. 15(a-c), respectively. The ergodic phase shows a sharper and longer ramp [Fig. 15(a)], whereas MBC phase shows a weaker and shorter ramp [Fig. 15(c)], indicating the presence of long-range correlations in the spectrum for both ergodic and MBC phases. In the MBL phase, no ramp regime is found indicating absence of spectral correlation [Fig. 15(b)] Prakash et al. (2021).
Appendix B Fock-space self energy and finite-size scaling
      Refer to caption(a)               Refer to caption(b)
Figure 16: System-size dependence of $\Delta_{t}$ for non-interacting EAAH model: (a) Log-log plots of $\Delta_{t}$ as a function of FS dimension $\mathcal{N}_{F}$ across the delocalization-localization transition with $\lambda$ for $\mu=0.5$. (b) Spectral dimension $D_{s}$ and pre-factor $A_{s}$ (inset), respectively, extracted from $\Delta_{t}=A_{s}\mathcal{N}_{F}^{-1+D_{s}}$ across the transition in (a). The vertical dashed line denotes the phase transition.
        Refer to caption(a)                Refer to caption(b)                 Refer to caption(c)                 Refer to caption(d)
Figure 17: System-size dependence of $\Delta_{t}$ and finite-size scaling for the interacting EAAH model: (a) Log-log plots of $\Delta_{t}$ as a function of FS dimension $\mathcal{N}_{F}$ across the ergodic-MBL transition with $\lambda$ for $\mu=0.5$. (b) Spectral fractal dimension $D_{s}$ and the pre-factor $A_{s}$, respectively, extracted from $\Delta_{t}=A_{s}\mathcal{N}_{F}^{-1+D_{s}}$ across the transition in (a). The vertical dashed line denotes the transition at $\lambda_{c}=3.2$. (c) Volumic scaling of $\Delta_{t}$ in ergodic phase where nonergodic volume $\Lambda$ shows a KT-like divergence approaching the ergodic-MBL transition. (d) Linear scaling of $\Delta_{t}$ in MBL phase where the correlation length $\xi_{s}$ shows power-law divergence approaching the transition. $D_{s}$ obtained from finite-size scaling analysis is shown in figure (b) with blue squares.
Here we first make an attempt to extract $A_{s}$ and $D_{s}$ from $\Delta_{t}=A_{s}\mathcal{N}_{F}^{-1+D_{s}}$ for delocalized-localized and ergodic-MBL transitions in non-interacting and interacting systems, respectively. $A_{s}$ and $D_{s}$, extracted through a linear fitting to the plots in Fig.Fig. 16(a) [also Fig. 5(a)], are shown in Fig. 16(b) across the delocalized-localized transition as a function of $\lambda$ for $\mu=0.5$ in the the non-interacting EAAH model. We find ($D_{s}\approx 1$, $A_{s}<1$) and ($0<D_{s}<1$, $A_{s}>1$) deep in the delocalized and localized phases, respectively.
Similarly, we extract $A_{s}$ and $D_{s}$ from a linear fitting of our numerical data in Fig. 17(a,b) for ergodic-MBL transition shown in Fig. 8(a). In the ergodic and MBL phases we find $(D_{s}\approx 1,A_{s}<1)$ and $(0<D_{s}<1,A_{s}>1)$, respectively. We draw the dashed vertical lines to denote the point $A_{s}=1$, which closely agree with the transition point $\lambda_{c}=3.2$ obtained from finite-size scaling analysis discussed below.
Similar to finite-size scaling of IPR discussed in the main text, we use the following scaling forms Sutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025),
$\ln\frac{\Delta_{t}}{\Delta_{c}}=\begin{cases}\mathcal{G}_{vol}\big(\frac{\mathcal{N}_{F}}{\Lambda_{s}}\big)&~~:~\text{Ergodic}\\ \mathcal{G}_{lin}\big(\frac{\ln\mathcal{N}_{F}}{\xi_{s}}\big)&~~:~\text{MBL}.\end{cases}$ (8)
Here $\Delta_{t}\sim\mathcal{N}_{F}^{-1+D_{s}}$ is the typical value of $\Delta_{I}$ in the thermodynamic limit, $\mathcal{N}_{F}\to\infty$; $\Delta_{c}\sim\mathcal{N}_{F}^{-1+D_{sc}}$ is its typical value at the critical point. According to the scaling ansatz, the scaling form inside the ergodic phase is volumic controlled by a ‘nonergodic volume’ $\Lambda_{s}$. In contrast,in the MBL phase, the scaling form is linear controlled by a length scale $\xi_{s}$. In the asymptotic limit of $\mathcal{N}_{F}\gg\Lambda_{s}$, the volumic scaling function in the ergodic phase follows $\mathcal{G}_{vol}\sim(1-D_{sc})\ln{(\mathcal{N}_{F}/\Lambda_{s})}$, where $\Delta_{t}\sim\Lambda_{s}^{(1-D_{sc})}$ Sutradhar et al. (2022). The asymptotic form of the linear scaling function for $\ln{\mathcal{N}_{F}}\gg\xi_{s}$ in the MBL phase follows $\mathcal{G}_{lin}\sim-(1-D_{sc})\ln{(\mathcal{N}_{F})}/\xi_{s}$, where $\xi_{s}=(1-D_{sc})/(D_{sc}-D_{s})$. The latter can be used to extract the spectral fractal dimension $D_{s}$ in the MBL phase. Figs. 10(c,d) show the volumic and linear scaling collpases in the ergodic and MBL phases, respectively, corresponding to phase transition discussed in Figs. 10(a,b). From the scaling analysis in the ergodic phase, we find $\mathcal{G}_{vol}(\mathcal{N}_{F}\gg\Lambda)\sim 0.35\ln{(\mathcal{N}_{F}/\Lambda_{s})}$ with $\lambda_{c}=3.2$, $D_{sc}\approx 0.65$ and $\Lambda_{s}\sim\exp{[B(\lambda_{c}-\lambda)^{-0.47}]}$ with $B\sim\mathcal{O}(1)$. In the MBL phase, $\mathcal{G}_{lin}(\ln{\mathcal{N}_{F}}\gg\xi_{s})\sim-0.35\ln{(\mathcal{N}_{F})}/\xi_{s}$ with $\xi_{s}\sim|\lambda-\lambda_{c}|^{-0.4}$. These exponents are similar to the ones reported in a recent study on ergodic-MBL transition in quasiperiodic AAH chain Ghosh et al. (2025) and close to the ones obtained from the analysis of IPR in this work. Although, in general, we find $D_{2c}<D_{sc}$, or even $D_{2}<D_{s}$, in the MBC and MBL phases (not shown).
Appendix C Distributions of the imaginary part of self-energy and inverse participation ratio
To gain further insights into the differences between MBC, and ergodic and MBL phases, we analyze the distributions of the imaginary part of FS self energy $\Delta_{I}$ over the FS sites in the middle slice of the FS lattice [Fig.4] and over realizations of disorder in all the phases of the noninteracting and interacting EAAH models. The distribution of $\Delta_{I}$ has been studied earlier in the ergodic, MBL and even in the NEE phases of other models Logan and Welsh (2019); Sutradhar et al. (2022); Roy et al. (2023).
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 18: Distributions of $\ln\Delta_{I}$: (a-c) Probability distributions $P(\ln\Delta_{I})$ in the ergodic $(\mu=0.5,\lambda=1.0)$, MBL $(\mu=0.5,\lambda=5.0)$ and MBC $(\mu=1.5,\lambda=1.0)$ phases, respectively. (d) $P(\ln\Delta_{I})$ in the delocalized $(\mu=0.5,\lambda=1.0)$ phase of the non-interacting EAAH model. Here solid lines are Gaussian fit to the data.
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 19: Distributions of IPRIPR: (a-c) Probability distributions of $IPR$ in the ergodic $(\mu=0.5,\lambda=1.0)$, MBL $(\mu=0.5,\lambda=5.0)$ and MBC $(\mu=1.5,\lambda=1.0)$ phases, respectively. (d) The same in the delocalized $(\mu=0.5,\lambda=1.0)$ phase of the non-interacting EAAH model.
In the ergodic phase, $\Delta_{I}$ has been shown to follow an almost log-normal (LN) distribution whereas in the non-ergodic phases it exhibits a broad distribution with a Levy-like power-law tail Logan and Welsh (2019); Roy et al. (2023); Sutradhar et al. (2022); Ghosh et al. (2025). In non-interacting systems with uncorrelated disorder on real-space lattices, both the delocalized and localized phases demonstrate LN distributions of real-space local self-energy  Schubert et al. (2010); Mirlin et al. (1996); Mirlin (2000). However, disorder on FS lattice originating from disorder, even if uncorrelated, of an underlying real-space lattice, becomes highly correlated. Thus, deviations from LN distributions have been obtained in the non-ergodic phases in earlier studies Roy et al. (2023); Ghosh et al. (2025).
We show distributions of $\ln\Delta_{I}$ in ergodic, MBL and MBC phases of interacting EAAH model in Figs. 18(a-c). For reference, we also show similar distributions of the FS self energy in the delocalized phase of the non-interacting EAAH model in Figs. 18(d). $P(\ln\Delta_{I})$ is a Gaussian for the LN distribution of $\Delta_{I}$. In the ergodic phase $P(\ln\Delta_{I})$, as shown in Fig. 18(a), is close to a Gaussian whose variance decreases with increasing $L$. In the MBL phase, $P(\ln\Delta_{I})$ is skewed and clearly deviates from a Gaussian. Here, the distribution broadens as $L$ increases [Fig. 18(b)], while developing a long tail. In the MBC phase, $P(\ln\Delta_{I})$ is closer to a Gaussian distribution, compared to in MBL. However, the variance of the distribution changes non-monotonically with $L$, as can be noticed in Fig. 18(c). In the non-interacting delocalized phase, shown in Fig. 18(d), $P(\ln\Delta_{I})$ is close to a Gaussian, but its variance does not change with $L$, unlike in the ergodic phase of the interacting system. The Anderson localized (AL) phase and critical phase of non-interacting systems show qualitatively similar $P(\ln\Delta_{I})$ as the MBL and MBC phases of the interacting EAAH model, respectively (not shown here). Thus, all the phases of the interacting system show some deviations from the LN distribution of $\Delta_{I}$, presumably due to the presence of correlated disorder on the FS lattice.
We also look into the distributions of FS IPR. In ref. Roy and Logan, 2021, the distributions of $D_{2}$ have been studied, assuming $A_{I}$ to be $\delta$-distributed, across the ergodic-MBL transition in a spin chain with random disorder. Here, we directly show the distributions of FS IPR, $P(IPR)$, in Figs. 19(a,b,c) for ergodic, MBL and MBC phases, respectively. The distribution in the ergodic phase [Fig. 19(a)] is quite distinct from those in the MBL [Fig. 19(b)] and MBC [Fig. 19(c)] phases, and even from the distribution in the non-interacting delocalized phase [Fig. 19(d)]. The distribution in the ergodic phase is rather skewed and exhibits longer tails compared to the MBC phase. The tail is absent in the MBL phase. The distribution is significantly less skewed in the delocalized phase of the non-interacting EAAH model. The distributions of IPR in the localized and critical phases (not shown) of the non-interacting EAAH model are similar to those in the MBL and MBC phases, respectively, of the interacting model.